# Implicit Bias of SignGD and Adam on Multiclass Separable Data

### v1

@@ 5 Implicit Bias of SignGD
We now use the descent property in (8) to lower bound the unnormalized margin. An intermediate result towards this is recognizing that sufficiently small loss $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ guarantees ${\bm{W}}$ separates the data (Lemma B.4). The descent property ensures that SignGD iterates will eventually achieve this loss threshold, thereby guaranteeing separability. The main result of this section, shows that eventually the iterates achieve separability with a substantial margin.
Lemma 5.2 (SignGD Unnormalized Margin).
@@
Assume $\tilde{t}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n},\forall t>\tilde{t}$. Then, the minimum unnormalized margin $\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}$ of iterates ${\bm{W}}_{t},t\geq\tilde{t}$ is lower bounded by ($\alpha_{s_{2}}=2Be^{2B\eta_{0}}$)
$\gamma\sum_{s=\tilde{t}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}-\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t-1}\eta_{s}^{2}.$ (9)
Proof.
By exponentiating the unnormalized margin:
(a) is because $\frac{\log(1+e^{-z})}{e^{-z}}\geq\log 2,z\geq 0$ with $z$ chosen to be $\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}$ for any $i\in[n]$; (b) is by some manipulations of (8) (details in App. C.2). ∎
SignGD Margin Convergence
@@
Proceeding from Eq. (9) requires showing convergence of the ratio $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}$. The two key ingredients are given in Lemma 4.3 (ii) and (iii). Lemma 4.3 (ii) suggests that it is sufficient to study the convergence of $\mathcal{L}({\bm{W}})$, which is captured in (8). However, to obtain an explicit rate via (8), we need to rewrite ${\mathcal{G}}({\bm{W}}_{t})$ in terms of $\mathcal{L}({\bm{W}}_{t})$. This is where Lemma 4.3 (iii) helps. Putting them together, we arrive at the following theorem.
Theorem 5.3.
Suppose that Ass. 3.1, 3.3, and 3.5 hold, then there exists $t_{s_{2}}=t_{s_{2}}(n,\gamma,B,{\bm{W}}_{0})$ such that the margin gap $\gamma-{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}/{\left\|{\bm{W}}_{t}\right\|_{\max}}$ of SignGD’s iterates for all $t>t_{s_{2}}$ is upper bounded by
$\mathcal{O}\Big{(}\frac{\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{s_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{s_{2}}-1}\eta_{s}+\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}}{\sum_{s=0}^{t-1}\eta_{s}}\Big{)}.$
Proof.
SignGD’s update rule (4) gives ${\left\|{\bm{W}}_{t}\right\|_{\max}}\leq{\left\|{\bm{W}}_{0}\right\|_{\max}}+\sum_{s=0}^{t-1}\eta_{s}$. We first find iteration $t_{s_{1}}$ via (8) such that $\mathcal{L}({\bm{W}}_{t+1})\leq\mathcal{L}({\bm{W}}_{t})-\frac{\eta_{t}\gamma}{2}{\mathcal{G}}({\bm{W}}_{t})$ for all $t\geq t_{s_{1}}$. Then, we find iteration $t_{s_{2}}>t_{s_{1}}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ for all $t\geq t_{s_{2}}$. By Lemma 4.3 (iii), this guarantees $\mathcal{L}({\bm{W}}_{t})\leq 2{\mathcal{G}}({\bm{W}}_{t})$. Substituting it to the above recursion on $\mathcal{L}({\bm{W}}_{t})$, we obtain $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}$. By Lemma 4.3 (ii), we further obtain $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\geq 1-e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}$. Combining this with (9) and the upper bound on ${\left\|{\bm{W}}_{t}\right\|_{\max}}$, the final result is proved (details in App. C.4). ∎
In the proof, $t_{s_{1}}$ can be set to $(4B^{2}e^{2B\eta_{0}}/{\gamma})^{1/a}$, and $t_{s_{2}}$ is chosen such that $t_{s_{2}}\leq\Theta(n^{\frac{1}{1-a}}+n^{\frac{1}{1-a}}\mathcal{L}({\bm{W}}_{0})^{\frac{1}{1-a}})$. Note that both are independent of the problem’s dimensionality, yielding the following convergence rates.
Corollary 5.4.
Set learning rate $\eta_{t}=\Theta(\frac{1}{t^{a}}),a\in(0,1]$. Under the setting of Theorem 5.3, the margin gap of SignGD iterates with $a=1/2$ reduces at rate $\mathcal{O}(\frac{\log t+n}{t^{1/2}})$. [^1] The rates for other values of $a$ can be found in Corollary C.5.
Remark 5.5.
These results generalize to NSD defined in (5). Concretely, Lemma 4.1, 4.2, and 4.3 (i) generalize to any $p$-norm (where margin in (2) is defined w.r.t. the same norm) together with the dual norm on the loss gradient. These results are summarized in Lemmas E.1 and E.3 in App. E. In terms of margin convergence of NSD, (Nacson et al., 2019) showed a rate of $\mathcal{O}(\frac{\log t}{t^{1/2}})$ in the binary setting, limited to the exponential loss. Compared to this, our results hold for the more practical setting of multilcass data and CE loss. The rate in Cor. 5.4 surpasses the max-norm margin convergence rate of Adam in the binary setting (Zhang et al., 2024); thus even in the binary setting it improves the latter and extends (Nacson et al., 2019)’s result to logistic loss. We elaborate on this gap in the next section.
6 Implicit Bias of Adam
Our analysis of Adam for multiclass data, while inspired by the binary setting of (Zhang et al., 2024), requires significant new technical insights to achieve the same convergence rate without class-dependent factors. While the proof follows similar high-level steps as SignGD above and leverages all the key properties of ${\mathcal{G}}({\bm{W}})$ identified in Sec. B, using ${\mathcal{G}}({\bm{W}})$ alone would lead to suboptimal bounds with an extra factor of $k$. Our crucial insight lies in a per-class decomposition of ${\mathcal{G}}({\bm{W}})$ as follows:
${\mathcal{G}}({\bm{W}})$ $=\sum_{c\in[k]}\frac{1}{n}\sum_{i\in[n],y_{i}=c}(1-s_{iyi})$
$=\sum_{c\in[k]}\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}s_{ic},$
where recall $s_{ic}\coloneqq\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})$. This decomposition motivates further defining the “per-class proxies”:
${\mathcal{G}}_{c}({\bm{W}})$ $\coloneqq\frac{1}{n}\sum_{i\in[n],y_{i}=c}(1-s_{iy_{i}})$
$\mathcal{Q}_{c}({\bm{W}})$ $\coloneqq\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}s_{ic}.$
Both terms play crucial roles in our per-class analysis, with this decomposition being key to avoiding the factor $k$ in our bounds. The remainder of this section highlights the technical challenges beyond those encountered in analyzing SignGD (full proof details are provided in Appendix D).
Adam Descent
To show descent of Adam via Taylor expansion (6), the first-order term involves an additional factor when introducing the dual norm of the gradient, i.e.,
$\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}_{t}\rangle\leq-\eta_{t}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}$
$\qquad+\eta_{t}\underbrace{\bigm{|}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigm{|}}_{\clubsuit}.$ (10)
Given that we bound the remaining terms in (6) using ${\mathcal{G}}({\bm{W}})$, it seems natural to bound the $\clubsuit$ term similarly. While one could attempt to relate the entries of both matrices $\bm{M}_{t}$ and ${\bm{V}}_{t}$ to ${\mathcal{G}}({\bm{W}}_{t})$, our entry-wise analysis of the $\clubsuit$ term (following (Zhang et al., 2024)’s approach in the binary setting) would inevitably introduce an extra factor of $k$. To avoid this, we instead relate the row vectors $\bm{M}_{t}[c,:]$ and ${\bm{V}}_{t}[c,:]$ to the class-specific functions ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$ for each $c\in[k]$. This class-wise approach requires careful accounting of interactions between the $k>2$ classes throughout our analysis. We first show how to bound $\bm{M}_{t}$.
Lemma 6.1.
@@ 6 Implicit Bias of Adam
Let $c\in[k]$. Under the setting of Theorem 6.3, there exists time $t_{0}$ such that for all $t\geq t_{0}$:
$|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})$ $\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|\leq$
$\alpha_{M}\eta_{t}\bigl{(}{\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})\bigr{)},$
where $j\in[d]$, and $\alpha_{M}:=B(1-\beta_{1})c_{2}$.
Here, we provide a quick proof sketch. By the Adam update rule 3a, we have for all $c\in[k]$ and $j\in[d]$: the following $|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|$ can be bounded by
$\sum_{\tau=0}^{t}(1-\beta_{1})\beta_{1}^{\tau}\underbrace{|\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|}_{\spadesuit}.$
By explicitly writing out the gradient and grouping terms we can show:
$\spadesuit$ $\leq B\frac{1}{n}\sum\nolimits_{i\in[n]}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|$
$=B\underbrace{\frac{1}{n}\sum\nolimits_{i\in[n]:y_{i}\neq c}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|}_{\spadesuit_{1}}$
$\quad\quad+B\underbrace{\frac{1}{n}\sum\nolimits_{i\in[n]:y_{i}=c}|\mathbb{S}_{y_{i}}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})|}_{\spadesuit_{2}}.$
In the above equality, we split the sum into two cases: samples where $y_{i}=c$ and the rest. Using the definitions of ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$, we manipulate and express the first ($\spadesuit_{1}$) and second ($\spadesuit_{2}$) terms via $\mathcal{Q}_{c}({\bm{W}})$ and ${\mathcal{G}}_{c}({\bm{W}})$ respectively. This leaves us to bound expressions of the form $|\mathbb{S}_{c}({\bm{v}}^{\prime})/\mathbb{S}_{c}({\bm{v}})-1|$ and $|(1-\mathbb{S}_{c}({\bm{v}}^{\prime}))/(1-\mathbb{S}_{c}({\bm{v}}))-1|$ respectively for the two terms. We address this technical challenge in Lemma D.10, showing how to bound such terms using $\|{\bm{v}}-{\bm{v}}^{\prime}\|_{\infty}$, which we can control for Adam. We now show how to bound ${\bm{V}}_{t}$.
Lemma 6.2.
Let $c\in[k]$. Under the setting of Theorem 6.3, there exists time $t_{0}$ such that for all $t\geq t_{0}$:
$\bigm{|}\sqrt{{\bm{V}}_{t}[c,j]}-\sqrt{(1-\beta_{2}^{t+1})}|$ $\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|\bigm{|}\leq$
$\alpha_{V}\sqrt{\eta_{t}}\bigl{(}{\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})\bigr{)},$
where $j\in[d]$, and $\alpha_{V}=B\sqrt{(1-\beta_{2})c_{2}}$.
By Adam’s update rule (3c), we have $\forall c\in[k],j\in[d]$ that $|{\bm{V}}_{t}[c,j]-(1-\beta_{2}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|$ can be bounded by
$\sum_{\tau=0}^{t}(1-\beta_{2})\beta_{2}^{\tau}\underbrace{|\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]^{2}-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|}_{\diamond}.$
For the $\diamond$ term, we compute $\nabla\mathcal{L}({\bm{W}})[c,j]^{2}$ and define the function $f_{c,i,p}({\bm{W}}):=(\delta_{cy_{i}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{i}))(\delta_{cy_{p}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{p}))$ to obtain (after some algebra):
$\diamond\leq\frac{B^{2}}{n^{2}}\sum\nolimits_{i\in[n]}\sum\nolimits_{p\in[n]}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|.$
Unlike the first moment $\bm{M}_{t}$, this summation involves pairs of sample indices $i$ and $p$ with label $c$ taking values equal to one of four cases: $y_{i},\text{non-}y_{i},y_{p},\text{non-}y_{p}$. Following our first-moment analysis strategy, we decompose the sum into these four components. The resulting bounds involve squared terms ${\mathcal{G}}_{c}({\bm{W}})^{2}$, $\mathcal{Q}_{c}({\bm{W}})^{2}$, and ${\mathcal{G}}_{c}({\bm{W}})\mathcal{Q}_{c}({\bm{W}})$ due to the quadratic softmax expressions in each component (see Lemma D.10 for bounding ratios of softmax quadratics).
With Lemmas 6.1 and 6.2 at hand, we can now bound the Adam-specific term $\clubsuit$ in Eq. (10) in terms of ${\mathcal{G}}({\bm{W}})$. The essence is to do the double sum (over $c\in[k]$ and $j\in[d]$) in $\clubsuit$ in two stages: first sum over $d$ for each $c\in[k]$ with the help of Lemma 6.1 and 6.2; then sum over $c$ with the recognition of ${\mathcal{G}}({\bm{W}})=\sum_{c\in[k]}{\mathcal{G}}_{c}({\bm{W}})=\sum_{c\in[k]}\mathcal{Q}_{c}({\bm{W}})$. The results are summarized in Lemma D.4. From this point on, following similar steps of SignGD, we can show the CE loss (eventually) monotonically decreases (see Lemma D.5).
Adam Implicit Bias
To establish Adam’s implicit bias, we follow the same strategy as SignGD: bound the margin and control weight growth. For the margin bound, we leverage our descent property from above (details in Lemma D.6). Unlike SignGD’s simple updates, Adam’s moment averaging makes bounding ${\left\|{\bm{W}}_{t}\right\|_{\max}}$ more challenging. Our solution again relies on the proxy ${\mathcal{G}}({\bm{W}})$, which we have shown approximates both loss and gradients well. Specifically, using this, we prove that the second moment ${\bm{V}}_{t}$ remains controlled when the loss is small, i.e. $\forall c\in[k],j\in[d]$:
${\bm{V}}_{t}[c,j]$ $\leq\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}+\alpha_{V}\eta_{t}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\leq 4B^{2}{\mathcal{G}}({\bm{W}}_{t})^{2}+\alpha_{V}\eta_{0}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\leq(4B^{2}+\alpha_{V}\eta_{0})\mathcal{L}({\bm{W}}_{t})^{2},$
where the penultimate and last inequalities are by Lemma 4.3 (i) and (ii), respectively. This suggests that with sufficiently small loss, all entries of ${\bm{V}}_{t}$ remain bounded by 1. Building on our bound for ${\bm{V}}_{t}$, we apply Zhang et al. (2024, Lemma A.4) (see also Xie & Li (2024, Lemma 4.2)), which connects ${\bm{W}}_{t}$ and $\log({\bm{V}}_{t})$ entry-wise. This allows us to bound ${\left\|{\bm{W}}_{t}\right\|_{\max}}$ using learning rate sums under Ass. 3.2 (details in Lemma D.7). With the convergence $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\rightarrow 1$ following as in SignGD, we arrive at our main theorem.
Theorem 6.3.
Suppose that Ass. 3.1-3.5 hold, and $\beta_{1}\leq\beta_{2}$. There exists $t_{a_{2}}=t_{a_{2}}(n,d,\gamma,B,{\bm{W}}_{0},\beta_{1},\beta_{2},\omega)$ such that the margin gap $\gamma-{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}/{\left\|{\bm{W}}_{t}\right\|_{\max}}$ of Adam’s iterates for for all $t>t_{a_{2}}$ is upper bounded by
$\mathcal{O}(\frac{\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{a_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{a_{2}}-1}\eta_{s}+d\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}^{3/2}}{\sum_{s=0}^{t-1}\eta_{s}}).$
The referenced iterations are $t_{a_{1}}=\Theta(d^{2/a})$ and $t_{a_{2}}\leq\Theta(n^{\frac{1}{1-a}}d^{2/a}+n^{\frac{1}{1-a}}\mathcal{L}({\bm{W}}_{0})^{\frac{1}{1-a}}+\log(1/\omega))$. Note that compared to the respective iterations for SignGD, $t_{a_{1}}$ depends on $d$, unlike $t_{s_{1}}$. These values the explicit rates below.
Corollary 6.4.
Set learning rate $\eta_{t}=\Theta(\frac{1}{t^{a}}),a\in(0,1]$. Under the setting of Theorem 6.3, the margin gap of Adam iterates with $a=2/3$ reduces at rate $\mathcal{O}(\frac{d\log(t)+nd+[\log(1/\omega)]^{1/3}}{t^{1/3}})$. [^2] The rates for other values of $a$ can be found in Corollary D.9.
Remark 6.5.
These rates exactly match those in the binary case of (Zhang et al., 2024) with logarithmic dependence on the initialization parameter $\omega$ (Ass. 3.2). This is only made possible through the fine-grained per-class bounding of the first and second moments using both ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$. Note that Lemma D.4 takes the same form as Zhang et al. (2024, Lemma A.4). However, without the tight per-class bound and the equivalent decomposition of ${\mathcal{G}}({\bm{W}})$ using either $\mathcal{Q}_{c}({\bm{W}})$ or ${\mathcal{G}}_{c}({\bm{W}})$, an extra factor of $k$ would appear. Interestingly, our rates for SignGD in Corollary 5.4 reveal a theoretical gap: Adam’s optimal choice $a=\frac{2}{3}$ yields $\mathcal{O}(\frac{d\log(t)+nd}{t^{1/3}})$ while SignGD achieves $\mathcal{O}(\frac{\log(t)+n}{t^{1/2}})$ with $a=\frac{1}{2}$. Despite achieving tightness w.r.t. class-dimension ($k$), this gap emerges from our entry-wise analysis of the $\clubsuit$ term in (10) across the feature dimension $(d)$ using scalar functions ${\mathcal{G}}_{c}({\bm{W}}),\mathcal{Q}_{c}({\bm{W}})$. Closing this theoretical gap–revealed through our SignGD analysis–that also appears in the binary case (Zhang et al., 2024), forms an important direction for future work.
Experiments

### v4

@@ 4 Implicit Bias of Normalized Steepest Descent
We now use the descent property in (9) to lower bound the unnormalized margin. An intermediate result towards this is recognizing that sufficiently small loss $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ guarantees $\bm{W}$} separates the data (Lemma 19 in App. C). The descent property ensures that NSD iterates will eventually achieve this loss threshold, thereby guaranteeing separability. The main result of this section, shows that eventually the iterates achieve separability with a substantial (unnormalized) margin.
Lemma 5 (NSD Unnormalized Margin).
@@
Assume there exists $\tilde{t}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n},\forall t>\tilde{t}$. Then, it holds that for all $t\geq\tilde{t}$ ($\alpha_{s_{2}}=2B^{2}e^{2B\eta_{0}}$)
$\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}\geq\gamma\sum_{s=\tilde{t}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}-\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t-1}\eta_{s}^{2}.$ (10)
NSD Margin Convergence
@@
Proceeding from Eq. (10) requires showing the convergence of the ratio $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}$. The two key ingredients are given in Lemma 3 (ii) and (iii). Lemma 3 (ii) suggests that it is sufficient to study the convergence of $\mathcal{L}({\bm{W}})$, which is captured in (9). However, to obtain an explicit rate via (9), we need to rewrite ${\mathcal{G}}({\bm{W}}_{t})$ in terms of $\mathcal{L}({\bm{W}}_{t})$. This is where Lemma 3 (iii) helps. Putting them together, we arrive at the following theorem (see Thm. 3 and Cor. 1 for details).
Theorem 1.
Suppose that Ass. 1, 2, and 4 hold. Set learning rate $\eta_{t}=\Theta(\frac{1}{t^{1/2}})$. The following holds for the margin gap of NSD’s iterates
$\gamma-\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{\lVert{\bm{W}}_{t}\rVert}\leq\mathcal{O}(\frac{\log t+n}{t^{1/2}}).$
Remark 1.
For margin convergence rates of NSD, nacson2019convergence showed a rate of $\mathcal{O}(\frac{\log t}{t^{1/2}})$ in the binary setting, limited to the entry-wise p-norms and the exponential loss. Compared to this, our results hold for the more practical setting of multiclass data and CE loss. To the best of our knowledge, this is the first non-asymptotic result on the implicit bias of spectral-GD for linear multiclass separable data, and it holds for other p-norms as well. Upon completion of this work, we became aware of an update on the arXiv version of tsilivis2024flavors, which includes an extension of their previous results to steepest descent w.r.t. the spectral norm. In comparison to ours, their gradient-flow analysis applies to homogeneous neural networks with the restriction of infinitesimal step-sizes. Moreover, it does not include normalization nor momentum (like Muon, which we analyze), and the convergence is (asymptotic) to a KKT point of a spectral-norm margin maximization problem.
5 Implicit Bias of Normalized Momentum Steepest Descent
In this section, we study the implicit bias of NMD algorithms (proof details in App E). Similar to Sec. 4, its updates are defined w.r.t. either the entry-wise or the Schatten norm $\lVert\cdot\rVert$. The analysis relies on the relationships in (7) and we show the same proxy function ${\mathcal{G}}({\bm{W}})$ naturally appears. Given that the NMD updates satisfy $\lVert\bm{\Delta}\rVert\leq 1$, the second-order term in (6) is bounded in the same way as NSD. The main difference is in bounding the first-order term as shown by the following lemma.
Lemma 6.
Let ${\bm{\Omega}}_{t}:=\bm{M}_{t}-\bm{\nabla}_{t}$. It holds for all $t\geq 0$ that
$\langle\bm{\nabla}_{t},{\bm{W}}_{t+1}-{\bm{W}}_{t}\rangle\leq 2\eta\lVert{\bm{\Omega}}_{t}\rVert_{*}-\eta\gamma{\mathcal{G}}({\bm{W}}_{t})\,.$
Given the relationships in (7) hold for any $p\geq 1$, we can bound the dual norm of $\lVert{\bm{\Omega}}_{t}\rVert_{*}$ via its sum norm (i.e. $\lVert{\bm{\Omega}}_{t}\rVert_{*}\leq{\left\|{\bm{\Omega}}\right\|_{\rm{sum}}}$). Given the goal is to bound all the terms in the Taylor expansion (6) via the proxy function ${\mathcal{G}}({\bm{W}}_{t})$, an natural next step is to bound $\left\|{\bm{\Omega}}_{t}\right\|_{\rm{sum}}$} using the same proxy function. To do this, we decompose the proxy function per-class-wise, and apply the per-class proxy functions to bound the entries of ${\bm{\Omega}}_{t}$ associated with their corresponding classes. Concretely, we write the function ${\mathcal{G}}({\bm{W}})$ in two equivalent ways: ${\mathcal{G}}({\bm{W}})=\sum_{c\in[k]}\frac{1}{n}\sum_{i\in[n],y_{i}=c}(1-s_{iyi})=\sum_{c\in[k]}\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}s_{ic}$, which motivate the following definitions of the per-class proxy functions:
${\mathcal{G}}_{c}({\bm{W}})\coloneqq\nicefrac{{1}}{{n}}\sum\nolimits_{i\in[n],y_{i}=c}(1-s_{iy_{i}}),\quad\text{and}\quad\mathcal{Q}_{c}({\bm{W}})\coloneqq\nicefrac{{1}}{{n}}\sum\nolimits_{i\in[n],y_{i}\neq c}s_{ic}.$
Next, we bound the entries in each row of ${\bm{\Omega}}_{t}$ (thus belonging to the same class) via the corresponding proxy functions ${\mathcal{G}}_{c}({\bm{W}}_{t})$ and $\mathcal{Q}_{c}({\bm{W}}_{t})$ to arrive at the following lemma. Its proof utilizes the nice properties of softmax map given in Lemma 15 in App. B.
Lemma 7.
@@ 5 Implicit Bias of Normalized Momentum Steepest Descent
Suppose that Ass. 1, 2, 3, and 4 hold. Let $c\in[k]$ and $j\in[d]$. There exists time $t_{0}$ such that for all $t\geq t_{0}$ and for $\alpha_{M}:=B(1-\beta_{1})c_{2}$:
$|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})$ $\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|\leq\alpha_{M}\eta_{t}\bigl({\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})\bigr)\,.$
Given the result in Lemma 7, we can show that $|{\bm{\Omega}}_{t}[c,j]|\leq\beta_{1}^{t+1}|\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|+\alpha_{M}\eta_{t}\mathcal{T}_{c}({\bm{W}}_{t})$, where $\mathcal{T}_{c}({\bm{W}}_{t})$ is defined to be $\mathcal{T}_{c}({\bm{W}}_{t}):={\mathcal{G}}({\bm{W}}_{t})+\mathcal{Q}({\bm{W}}_{t})$. Then, we sum over indices $c\in[k]$ and $j\in[d]$ and apply ${\left\|\bm{\nabla}\right\|_{\rm{sum}}}\leq 2B\cdot{\mathcal{G}}({\bm{W}})$ (from Lemma 3 (i)) to obtain:
Lemma 8.
It holds for all $t\geq 0$ that ${\left\|{\bm{\Omega}}_{t}\right\|_{\rm{sum}}}\leq 2B\beta_{1}^{t/2}{\mathcal{G}}({\bm{W}}_{t})+2\alpha_{M}d\eta_{t}{\mathcal{G}}({\bm{W}}_{t})$.
This completes the bound on the first-order term for NMD algorithms via the proxy ${\mathcal{G}}({\bm{W}})$. The rest proof follows similar steps as NSD. We note that without the above per-class decomposition, an extra $k$-factor would appear in the second term of the bound on $\left\|{\bm{\Omega}}_{t}\right\|_{\rm{sum}}$} (and thus also show up in the final rate). We state the main theorem for NMD algorithms.
Theorem 2.
Under the setting of Lem. 7, the margin gap of NMD with $\eta_{t}=\Theta(\frac{1}{t^{1/2}})$ is $O(\frac{d\log t+dn}{t^{1/2}})$.
Remark 2.
wang2022does studied implicit bias of un-normalized GD with momentum, and showed its iterates converge asymptotically to the max 2-norm margin solution. In contrast, our rates are non-asymptotic and cover a much wider family of algorithms converging to non-Euclidean geometric margins (w.r.t. entry-wise/Schatten norms). Note the convergence rate of NMD matches that of NSD (Thm. 1) up to a factor of $d$. It could be interesting to remove this dependence in a future work.
Implicit Bias of Adam
Finally, observing that Adam [kingma2014adam] (without the stability constant, i.e., eqns (33a), (33b), and (33c) in App. G) shares the same form of momentum as NMD and the (entry-wise) updates are bounded by some constant as shown in zhang2024adam and xie2024implicit. Thus, our analysis extends to Adam. Concretely, a similar proof strategy can be adapted once a bound on the second gradient moment via the proxy function is established (Lemma 35). In App. G, we prove a $O(\frac{d\log(t)+nd}{t^{1/3}})$ max-norm margin convergence rate for Adam (details in Thm. 5 and Cor. 2).
6 Experiments
Synthetic Experiments